The difference of two numbers is 48. The second is 8 less than 5 times the first. What are the two numbers?
Answer by wgunther(43)

Let the first number be a and the second be b. so a-b=48, b+8=5a. Solve this system of equations. you can add them, so a-b+b+8=48+5a. so a+8=48+5a. so 4a=-40. so a=-10. b=-10-48=-58